### v1

# Planar near-field measurements of specular and diffuse reflection of millimeter-wave absorbers

Planar near-field measurements of specular and diffuse reflection of millimeter-wave absorbers
††journal: opticajournal††articletype: Research ArticleMitigating the far sidelobes of a wide field-of-view telescope is one of the critical issues for polarization observation of the cosmic microwave background. Since even small reflections of stray light at the millimeter-wave absorbers inside the telescope may create nonnegligible far sidelobes, we have developed a method to measure the reflectance of millimeter-wave absorbers, including diffuse reflections. By applying the planar near-field measurement method to the absorbers, we have enabled two-dimensional diffuse-reflection measurements, in addition to characterizing specular reflection. We have measured the reflectance of five samples (TK RAM Large and Small Tiles and Eccosorb AN-72, HR-10, and LS-22) at two angles of incidence in the frequency range from 70 GHz to 110 GHz. Compared with conventional horn-to-horn measurements, we obtained a consistent specular reflectance with a higher precision, less affected by standing waves. We have demonstrated that the angular response and diffuse-to-specular reflectance ratio differ among various materials. The measurements also imply that some absorbers may affect the polarization direction when reflecting the incident waves.
1 Introduction
The primordial gravitational waves due to inflation are believed to have left a characteristic polarization pattern called modes in the cosmic microwave background (CMB)[1]. The -mode polarization is expected to peak at large angular scales of several degrees or more, and the stray light contamination due to emission from the galactic plane entering the telescope is expected to be one of the largest systematic errors[2, 3, 4]. Stray light can enter the telescope through the far sidelobes. Next-generation CMB telescopes such as LiteBIRD need to know sidelobe to an accuracy of -56 dB[5].
To achieve such a requirement, it is essential to characterize reflections from millimeter-wave absorbers[6], for example, at the baffling walls, in addition to minimizing stray light in the optical design optimization[7]. In recent years, novel millimeter-wave absorbers have been reported, such as metamaterial absorbers[8, 9], 3D printed absorbers[10, 11, 12], and new carbon-based absorbers[13, 14].
The reflectance properties of millimeter-wave absorbers in free space have been made using two horns and focusing optics, such as mirrors[9], in such a way that the attenuation of the coupling constant between the horns is measured (horn-to-horn method). Since stray light enters the absorber from various angles inside the telescope, the specular reflectance is measured at multiple angles of incidence[15, 16]. Also, by rotating one of the horn antennas around the absorbers, one-dimensional distributions of diffuse reflection have been measured[17, 18, 9]. However, if the scattered waves from the absorber are only partially coupled to the receiving horn, the measured reflectance is underestimated.
In this paper, we report a new method to measure the reflectance of millimeter-wave absorbers. Applying the planar near-field antenna measurement method[19, 20], we measure the amplitude and phase of the reflected waves from absorbers on a plane. By applying plane wave expansion to the measurements, we characterize the two-dimensional angular patterns of diffuse reflection, from which the specular reflectance is also evaluated.
We characterized the specular and diffuse reflectance for five absorbers (AN-72, HR-10, LS-22, TK RAM Large, and TK RAM Small) in the 70 – 110 GHz range. Since the reflectance characteristics of the absorber depend on the polarization direction[21, 17], we performed measurements for both incident polarizations (S-pol and P-pol). We set the polarization direction of the probe not only parallel but also perpendicular to the polarization direction of the feed horn.
The structure of this paper is as follows. Section 2 describes the principles of the new measurement method proposed in this paper. Section 3 describes the measurement setup constructed to demonstrate the measurement method. Section 4 reports the measurement results of millimeter-wave absorbers and discusses the specular and diffuse reflectance. Section 5 states the conclusion.
2 Reflection measurement method
2.1 Horn-to-horn method
Conventional millimeter-wave absorber reflection measurements have been made using two horns and focusing optics, such as mirrors[9] in such a way that the attenuation of the coupling constant between the horns is measured (horn-to-horn method; Fig. 1(a))[16, 9]. This method has been extended to one-dimensional diffuse reflection characterization by changing the angle of the receiving horn ( in Fig. 1(a))[17, 18]. However, if the scattered waves from the absorber are only partially coupled to the receiving horn, the measured reflectance is underestimated. It has been reported that the measured reflectance depends on the distance between the absorber and the receiving horn[17].
2.2 Planar near-field method
In order to accurately characterize absorber reflection, including diffuse reflection, we have applied the planar near-field measurement method[19, 22] to the reflection measurement of millimeter-wave absorbers. Figure 1(b) shows a schematic diagram of the proposed method. In this method, we illuminate the test absorber with a feed horn and measure the waves reflected from the test absorber by moving the receiver with the probe on a two-dimensional plane. The measured amplitude and phase distributions are decomposed into plane-wave components with different angular directions by plane wave expansion (two-dimensional Fourier transform) as[23, 24]
Here, is the complex electric field reflected by the test absorber; denote the spatial frequency in the plane, which relate the angle as and with being the wave number, wavelength and speed of light, respectively. is the azimuth angle from the -axis.
The planar near-field measurement has three advantages. First, by measuring the reflected wave of the absorber, we can evaluate the reflectance characteristics in more detail. Compared with the conventional horn-to-horn method, the information obtained from the two-dimensional distribution of reflected waves, including diffuse reflections, is much greater. Second, since there is no need to know the reflection point of the absorber, it is less susceptible to alignment adjustments in optical measurements. Third, compared with horn-to-horn measurements, measurements with a small-aperture probe enable more accurate measurements less affected by standing waves.
We evaluate the specular and total reflectance of the absorbers by integrating the measured two-dimensional angular reflectance into different integration ranges. The former is obtained by the integration of the reflected waves in the Half Width at Half Maximum (HWHM) of the beams of the feed horn,
which is considered to be similar to the conventional horn-to-horn measurement. The latter is obtained by the integration of the entire two-dimensional angular reflectance (),
Here, the angular response is calculated for each frequency using Eq. (1).
3 Experimental verification
3.1 Experimental setup
Figure 2 shows the measurement setup developed in this study. We measured the amplitude and phase of the electric field reflected by the test absorbers using a vector network analyzer (VNA: Keysight Technologies N5222B) and frequency extenders (Virginia Diodes WR10VNAX). Based on the reciprocity theorem, the coupling between the two horns is equivalently measured when the VNA transmitter and receiver are reversed[25]. For the convenience of the setup, we connected the probe and lens horns to the transmitter and receiver, respectively, in this experiment.
The measurements were performed with a 1 GHz interval over a frequency range of 70 GHz – 110 GHz. Signals from 11.6 to 18.3 GHz were output from the VNA and converted to six times higher frequencies. Then, the 70 – 110 GHz signal is output from the probe. We measured the reflected waves from the millimeter-wave absorber by moving the probe in 1.5 mm increments in a 75 mm 75 mm square area. The distance between the probe and test absorbers is set to 60 mm, considering mechanical clearance. This configuration measures the specular and diffuse reflection in . The reflected waves are received by a conical horn (aperture diameter 23.0 mm, opening angle 10.0∘) with a polymethyl pentene (TPX) lens (diameter 25.4 mm, focal length 50 mm) placed in the aperture. The distance between the lens horn and test absorbers is 50 mm.
The millimeter-wave absorber under test is placed on two rotating and one-axis stages. Since the movable range of the measurement stage is finite, in addition to the angle of incidence of the lens horn (AoI in Fig. 1(b)), the angle of the absorber relative to the measurement stage ( in Fig. 1(b)) was also rotated so that the measurement plane and the signal optical path are vertical. In addition, using a one-axis stage, we aligned the surface of the absorbers at a common plane for all the measurements. Considering the situation in telescopes, we measured the absorbers fixed on a 5 mm thick aluminum plate. The absorbers and the aluminum plate are 200 mm 200 mm in size. An acrylic frame held the absorber in place, leaving no space between it and the aluminum plate.
Table 1 summarizes the measurement parameters for this experiment. We measure the reflectance for two polarizations; the polarization direction of the lens horn is set parallel to the and axes in Fig. 1 (S- and P-pol, respectively). In addition, the polarization direction of the probe is set not only parallel but also perpendicular to that of the lens horn.
Frequency range | 70 - 110 GHz |
Frequency step | 1 GHz |
Angle of incidence (AoI) | 30 deg, 45 deg |
Co-polarization of the lens horn | S-pol, P-pol |
Probe polarization | ss, pp (parallel to the lens horn) |
sp, ps (perpendicular to the lens horn) |
3.2 Millimeter-wave absorbers
Samples | Material | Form | Structure | Thickness | Mass | Ref. |
[mm] | [kg/m2] | |||||
AN-72 | polyurethane | sheet | multi-layer | 6 | 1.0 | [26] |
HR-10 | polyurethane | sheet | single-layer | 10 | 0.6 | [27] |
LS-22 | polyurethane | sheet | single-layer | 9.8 | 0.5 | [28] |
TK-L | polypropylene | tile | pyramid | 14.9 | 10.6 | [29] |
TK-S | polypropylene | tile | pyramid | 7.6 | 5.9 | [29] |
Table 2 summarizes the test millimeter-wave absorbers measured in this experiment. Eccosorb AN-72[26], HR-10[27], and LS-22[28] are commercially available millimeter-wave absorbers by Laird. They are made of polyurethane-based foams with different structures, which are expected to have different diffuse reflection characteristics: LS-22 has a single-layer structure; AN-72 has a multi-layer structure; HR-10 has a single-layer structure and contains carbon powder. LS-22 is available in several thicknesses, and we measured a 3/8-inch one. Eccosorb AN-72 and Eccosorb HR-10 are used in CMB ground-based experiments[30, 31, 32]. TK RAM is a radar absorber developed by Thomas Keating for the terahertz and millimeter-wave bands, and there are two types of TK RAM[29]. TK RAM Small tile (abbreviated as TK-S) is a small tile of 25 mm 25 mm and is recommended for 200 – 600 GHz. TK RAM Large tile (abbreviated as TK-L) is a large tile of 100 mm 100 mm and recommended for 50 – 200 GHz. TK-L and TK-S have different-sized surface structures, 5.0 mm, and 2.5 mm, respectively.
3.3 Reference measurement with an aluminum plate
Assuming that the aluminum plate without the absorber is a perfect conductor[33], we measured it for intensity calibration.
Figure 3 shows reflection measurements of the reference aluminum plate. The upper panels show the angular response of the electric field parallel and perpendicular to the co-polarization of the S and P-polarized lens horn at 90 GHz. The angular response are calculated using Eq. (1), and is normalized by the peak value of . The HWHM of the lens horn was measured to be 4.34 deg at 90 GHz, consistent with the prediction by a Gaussian-beam approximation[34]. The cross-polarization level of the lens horn was measured to be dB at 90 GHz. The lower panel shows the profile in a cut-plane. We confirm that this setup can be used to measure for two polarizations since the patterns measured with both polarizations are consistent at the dB level over a range of .
4 Results and discussion
4.1 Planar near-field reflection measurements
We have measured the reflectance of the millimeter-wave absorbers using the planar near-field measurement with the setup described in Sec. 3.1. Figure 4 shows the amplitude and phase distributions of the measured millimeter-wave absorber reflections with ss polarization (S-polarized lens horn; probe polarization parallel to the lens horn) at 90 GHz for AoI = 30∘. The amplitude distributions are normalized by the peak value of the amplitude distribution of the reference aluminum plate for each polarization. Compared with AN-72, HR-10 sample has a more spread pattern.
4.2 Angular responses and reflectance of the absorbers
Figure 5 shows the angular responses of the reflected waves from the absorbers at 90 GHz for AoI = 30∘. The results are normalized by the peak value of the reference aluminum plate for each polarization. AN-72 has an almost isotropic attenuation from the reference aluminum plate pattern, while HR-10 has a spread pattern. LS-22 has similar patterns to AN-72, but the difference in S-pol and P-pol is smaller than that of AN-72. TK-S also shows a pronounced diffuse reflection pattern, indicating an upward tilt. This may be caused by the surface structure and errors in attaching TK-S to the reference aluminum plate.
Figure 6 shows the angular responses at 90 GHz for AoI = 45∘. It is expected that a small but direct signal from the transmitter to the receiver is present. Such a direct path between the lens horn and probe was also measured without placing a millimeter-wave absorber. As the incident angle increases, the direct path becomes more extensive, and stripe patterns due to interference with the direct path are observed more apparently than those of S-pol in the case of P-pol. This is because the sidelobe level of the lens horn is higher in the polarization direction. This may be improved in future measurements by using horns with small sidelobes, such as a corrugated horn.
Figure 7 is the cross-sectional profiles of Fig. 5 for , and . While AN-72 and LS-22 have a constant attenuation from the reference aluminum plate pattern, the intensity of the reflected waves from HR-10 ranges between -40 and -20 dB regardless of the angle . Notably, in the angular range of 20 30∘, the level of the reflected waves from HR-10 is higher than that of the reference aluminum plate.
Figure 8 compares the specular and total reflectance of the absorbers in the S- and P-polarization, calculated using Eqs. (2) and (3). Here, in Eq. (3) is set to 30 degrees. The gray line shows of the direct path from the probe to the lens horn measured without placing a millimeter-wave absorber. The effects of direct path is about 10 dB larger in P-pol than in S-pol. The specular reflectance of AN-72 is about -12 dB for S-pol and -17 dB for P-pol, which is about 5 dB different for the two polarizations. On the other hand, the specular reflectance of HR-10 is about -20 dB for both polarizations, which is smaller than that of AN-72. In general, dielectric materials have more significant reflectance for the P-pol than the S-pol for oblique incident wave[21]. The measurements of AN-72 show a consistent tendency. By contrast, the reflectance of LS-22 shows little difference between the two polarizations. The reflectance of TK-L was about -40 dB for both polarizations, the lowest among the five absorbers measured in this experiment.
The specular reflectance (green line) of HR-10 for the S-polarization is about -20 dB, while it rises to about -15 dB for total reflectance (red line). This shows that HR-10 has larger diffuse reflection than other absorber, consistent with the isotropic spreading feature of angular patterns in Fig. 5. The total reflectance of TK-S is also higher than the specular reflectance by around 5 dB. For other absorbers, the two reflectances have little difference, indicating that diffuse reflection is small.
We also compared the measured specular reflectance discussed in Sec. 2.1 with conventional horn-to-horn measurements. We performed the horn-to-horn measurements using the same experimental setup in Fig. 1[Sec. 3.1] by replacing the probe with the lens horn. The two lens horns for the transmitter and receiver are identical. The results demonstrate that the planar near-field measurement and the horn-to-horn reflection measurement agree within a few dB of accuracy and that the near-field measurements are less affected by standing waves (Fig. 9).
4.3 Electric field measurements perpendicular to the co-polarization of the lens horn
Using the planar near-field measurement method, we also measured the electric field perpendicular to the co-polarization of the lens horn. Figure 10 shows the angular responses of the electric field perpendicular to the co-polarization of the lens horn at 90 GHz for AoI = 30∘. The patterns of AN-72, LS-22, and TK-S measured with the S- and P-polarized lens horn (labeled as “sp” and “ps” respectively) are similar to those of the reference aluminum plate (“sp” and “ps” in Fig. 3) with attenuated levels, showing a four-divided pattern. This indicates that the patterns are determined mainly by the cross-polarization of the lens horn. On the other hand, the patterns of HR-10 and TK-L do not resemble that of the reference aluminum plate.
Such trends are also apparent in total reflectance defined by Eq. (3), as summarized in Fig. 11. For AN-72, LS-22, and TK-S, the total reflectance measured with the S- and P-polarized lens horn ( and respectively) are mostly consistent with and . Here, and indicate the cross-polarization components of the lens horn, which are obtained from the reference aluminum plate measurements as,
However, the total reflectance and of HR-10 and TK-L are dB larger than and . The measurements imply that HR-10 and TK-L may affect the polarization direction when reflecting the incident waves.
5 Conclusion
By applying the planar near-field antenna measurement method, we have developed a method to measure the reflectance of millimeter-wave absorbers, including the two-dimensional distribution of diffuse reflections. We have measured five absorbers (TK-L, TK-S, and Eccosorb AN-72, HR-10, LS-22) at and angles of incidence for two polarizations (S-pol and P-pol) in the frequency range from 70 GHz to 110 GHz. From these measurements, we evaluate the specular and total reflectance of each absorber. The obtained specular reflectance is less affected by standing waves and agrees with the conventional measurement. The specular reflectance of AN-72 and HR-10 are higher for S-pol, while that of LS-22, TK-L, and TK-S does not depend on incident polarization. We also evaluate the total reflectance, including diffuse reflection. The diffuse-to-specular reflectance ratio is significantly higher for HR-10 and TK-S compared with AN-72 and LS-22. We also measured the electric field perpendicular to the co-polarization of the lens horn. The measurements imply that HR-10 and TK-L may affect the polarization direction when reflecting the incident waves.
While the conventional horn-to-horn method measures the power coupling efficiency between the two horns, the present method allows the detailed near-field mapping of the wave reflected from the absorber. Thus, the latter can investigate the nature of the reflection more extensively. We have demonstrated in this paper that the diffuse-to-specular reflectance ratio is different among various materials. It is also a merit of the new method that the measurement is not sensitive to the alignment of the absorber in the setup optics. Practically, it will help us conduct similar reflection measurements in cryogenic conditions. The method is expected to contribute to the development of millimeter-wave absorbers by achieving accurate measurements, including diffuse reflections, and to make a better design against the stray light in CMB observation telescopes.
Funding This work was supported by MEXT/JSPS KAKENHI Grant Numbers 23K25889, JP23K17309, JP23K25889 and JP24K17078 and by JSPS Core-to-core Program Number JPJSCCA20200003.
Acknowledgments We are grateful to Luca Lamagna, Andrea Occhiuzzi, Haruaki Hirose and Takuro Fujino for their helpful discussions on the measurements. We are also grateful to Oshima Prototype Engineering Co. for fabricating the horn antennas used in the measurements.
Disclosures The authors declare no conflicts of interest.
Data Availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.